zarkam21 (zarkam21):

Solve the system of equations. Enter your answer as an ordered pair. 4x+6y=62 -6x+6y=-18

OpenStudy (anonymous):

lets use elimination method

Mehek (mehek14):

eliminate y by subtracting

Mehek (mehek14):

4x + 6y = 62 - (-6x) + 6y = -18 ----------------- 10x + 0 = 80 x = 8 now substitute

zarkam21 (zarkam21):

(8,-6)

Mehek (mehek14):

no not -6

Mehek (mehek14):

4 (8) + 6y = 62 32 + 6y = 62 6y = 30 y = 5

zarkam21 (zarkam21):

So (8,5) ?

Mehek (mehek14):

yea